If 20 g steam initially at 100 degree C is added to 60 g of ice initially at 0 degree C, then find the final

equilibrium temperature of the mixture.


1
Expert's answer
2021-08-16T17:45:07-0400

Gives

"m_1=20g\\m_2=60g\\T_1=100°c\\T_2=0°c"

"T=\frac{m_1T_1+m_2T_2}{m_1+m_2}"

"T=\frac{20\times100+60\times0}{20+60}=\frac{2000}{80}=25°c"